OpenStudy (nottim):

Solve x for 0=xe^(2x)

OpenStudy (deoxna):

The only time f(x)=0 is when x=0, because there is no value of x that will make e^2x=0.

OpenStudy (nottim):

I was thinking of using ln, but then i remember ln0 not right

OpenStudy (ranga):

If A * B = 0 either A is 0 or B is 0.

OpenStudy (anonymous):

\[e ^{2x}>0 (always)\] so only solution is x=0

OpenStudy (deoxna):

@NotTim You are right. ln is not defined for x=0.

OpenStudy (nottim):

Somewhat unrelated, but the derivative is 2x(x)e^(2x)+e^2x, correct?

OpenStudy (deoxna):

Sorry, that should be ln is not defined for\[x \le 0\]

OpenStudy (ranga):

2xe^(2x)+e^2x = e^(2x)(2x + 1)